Hi
bob had a total of 720 apples and oranges. 70% of them were oranges. He bought some more applesso that tthe percentage of apples increased to 40% of the total number of fruits. How many apples did he buy.
thanks

Answer by htmentor(1343)

Initially oranges make up 70% of the total. So the number of oranges = 0.7*720 = 504
Thus the number of apples = 720 - 504 = 216
Let a = the additional number of apples bought
Then we have 216 + a = 0.4(720 + a), since the new amount of apples is 40% of the new total
Solving for a gives a = 120
So the total number of apples bought is 336